# PPA-Game: Characterizing and Learning Competitive Dynamics Among Online Content Creators

### v1

@@ 3.3.2 Existence of PNE in Specific Scenarios
[…]
@@
The proof is provided in Appendix C.1. Theorem 3.1 demonstrates scenarios with a long-tailed distribution of resource payoffs. In particular, Equation (3) specifies that the peak payoff among resources substantially surpass that of other resources. For instance, entertainment-related content on TikTok garners substantially more views than other topics (Statista, 2023).
@@
The proof is provided in Appendix C.2. Theorem 3.2 demonstrated that the PNE exists in several degraded cases of the PPA-Game in Definition 3.2. Specifically, Condition (1) corresponds to scenarios where the weights reflect the inherent influence of different players. For instance, in online recommender platforms, certain creators, termed as ‘influencers,’ have greater impact on all topics owing to their expertise, stature, or authority. We also note that Condition (1) exhibits a special case of the weighted congestion game (Bhawalkar et al., 2014). However, PNE may not exist in weighted congestion games while we prove the existence in this specific scenario. Condition (2) encompasses cases where players are homogeneous and have the same influence on each topic. This scenario aligns with the one addressed in (Xu et al., 2023). Condition (3) considers a homogeneous resource scenario wherein all resources have equivalent payoffs. Such scenarios include that users compete for processing resources in modern data centers while each resource has the same processing power (Benblidia et al., 2021).
@@ 3.3.3 Non-uniqueness and Inefficiency of PNEs
Given the existence for such inefficient PNEs, it becomes preferable for players to converge to an optimal PNE, thereby maximizing the overall welfare.
@@ 4.1 Multi-Player Multi-Armed Bandits Framework
When multiple players select the same arm, the reward is allocated according to the proportional payoff rule specified in Equation (1). Specifically, let $R_{j}(t)$ represent the reward received by player $j$ at round $t$, given by:
$R_{j}(t)=X_{k}(t)\cdot\frac{w_{j,k}}{\sum_{j^{\prime}=1}^{N}\mathbb{I}[\pi_{j^{\prime}}(t)=k]w_{j^{\prime},k}},\,\text{where}\quad k=\pi_{j}(t).$ (4)
@@ 4.2 Evaluation metrics
Notably, $\operatorname{Reg}_{j}(T)$ is a random variable and the randomness come from the actions $\{\bm{\pi}(t)\}$, which relies on historical random observations $\{X_{k}(t)\}$.
@@ 4.3 Algorithm
Inspired by (Bistritz and Leshem, 2018, 2021; Pradelski and Young, 2012), we introduce a novel online learning algorithm designed to enable players to both learn the game’s parameters and optimize their individual rewards concurrently. The algorithm’s framework is shown in Figure 1 and pseudo-code is outlined in Algorithm 1.
@@ 4.3.2 Learning and Exploitation Phase
In this subphase, we adopt the framework proposed in (Pradelski and Young, 2012) to identify the most efficient Nash equilibrium. Specifically, each player maintains a status tuple $(m_{j},a_{j},u_{j})$ at each round $t$ during this phase. Here, $m_{j}$ can be one of four moods: discontent ($D$), content ($C$), watchful ($C^{-}$), and hopeful ($C^{+}$). Additionally, $a_{j}$ denotes the arm currently being played by player $j$, and $u_{j}$ represents the utility player $j$ gains when playing arm $a_{j}$.
Intuitively, the concept of mood shows a player’s attitude towards their current strategy. A discontent player is uncertain about which arm is the most profitable, necessitating further exploration. Conversely, a content player believes they are already following the most efficient PNE. A watchful player is skeptical about their current strategy’s efficacy and may transition to a discontent mood. Finally, a hopeful player believes that their current strategy will yield higher-than-expected returns, and thus continues to adhere to it.
At each round $t$, players pull arms in accordance with Rule 4.1 based on their mood from the previous round.
@@
The chosen arm $\pi_{j}(t)$ is chosen based on the mood at the last round.
@@
Rule 4.1 captures the essence of the different moods. Specifically, discontent players explore the arm space uniformly, while content players predominantly stick to their current strategy but also explore other strategies with low probability. Watchful and hopeful players require additional rounds to ascertain the efficacy of their current strategy, so they maintain their current choice of arm.
@@
Define the functions $F(u)$ and $G(u)$ as follows:
@@
    $(m_{j},a_{j},u_{j})\leftarrow\begin{cases}(C^{+},a_{j},u_{j})&\text{if }u^{\prime}_{j}>u_{j},\\ (C,a_{j},u_{j})&\text{if }u^{\prime}_{j}=u_{j},\\ (C^{-},a_{j},u_{j})&\text{if }u^{\prime}_{j}<u_{j}.\end{cases}$ (10)
@@
    $(m_{j},a_{j},u_{j})\leftarrow\begin{cases}(C^{+},a_{j},u_{j})&\text{if }u^{\prime}_{j}>u_{j},\\ (C,a_{j},u_{j})&\text{if }u^{\prime}_{j}=u_{j},\\ (D,a_{j},u_{j})&\text{if }u^{\prime}_{j}<u_{j}.\end{cases}$ (11)
@@
    $(m_{j},a_{j},u_{j})\leftarrow\begin{cases}(C,a_{j},u^{\prime}_{j})&\text{if }u^{\prime}_{j}>u_{j},\\ (C,a_{j},u_{j})&\text{if }u^{\prime}_{j}=u_{j},\\ (C^{-},a_{j},u_{j})&\text{if }u^{\prime}_{j}<u_{j}.\end{cases}$ (12)
@@
Rule 4.2 is designed to reflect the intuitions behind the various moods. For discontent players, the status transitions to content if the randomly chosen strategy yields high utility. Content players, on the other hand, adjust their status based on the observed utility at the current round if they have not explored other strategies; otherwise, they transition to a different strategy based on utility. Watchful and hopeful players adjust their status based on the observed utility as well.
In line with existing work (Pradelski and Young, 2012; Young, 2009; Marden et al., 2014), Rule 4.1 and 4.2 constitute a trial-and-error learning procedure, akin to common learning routines in human behavior (Young, 2009).
Additionally, during the Learning PNE Subphase, each player maintains a counter $Q_{j,k}$ for each arm $k\in[K]$. This counter records the number of times that arm $k$ has been chosen while the player is in a Content mood.
@@ 4.4.1 Quantifying the Challenge of Identifying the Most Efficient PNE
$\delta^{\text{NE}}$ $=\min\big{\{}\mathcal{U}_{j}(\bm{\pi})-\mathcal{U}_{j}(\pi^{\prime},\bm{\pi}_{-j}):\bm{\pi}\in A^{\text{NE}},j\in[N],\pi^{\prime}\in[K],\pi^{\prime}\neq\pi_{j}\big{\}}.$ (13)
@@
Set $N=3$ and $K=2$. Here we suppose that each player has the same weights on the resources (as per Condition (2) in Theorem 3.2). When arms’ rewards are $[1,0.7]$ and players’ weights are $[1,0.8,0.4]$, we have $0.0260.052=\delta^{\text{NE}}>\delta^{\text{NoNE}}=$ 0.026. By contrast, when arms’ rewards are $[1,0.6]$ and players’ weights are $[1,2/3,4/9]$, we have $0.0680.040=\delta^{\text{NE}}<\delta^{\text{NoNE}}=$ 0.068. Details of the two cases can refer to Tables 1 and 2 in Appendix B.
Let $\Delta$ be the largest discrepancy between the estimated and true rewards $\Delta=\max_{j,k}\left|\hat{\mu}_{j,k}-\mu_{k}\right|$. We can have the following theorem.
@@
Suppose $\Delta<\delta/(4K)$, $\Gamma\leq\delta/(4N)$, and the PNE exists for the game $\mathcal{G}$ in Definition 3.2. Then the PNE exists for the game $\mathcal{G}^{\prime}$ (Equation (7)). In addition, the most efficient PNE in the game $\mathcal{G}^{\prime}$ is unique with probability 1 and the equilibrium is a most efficient PNE in the original game $\mathcal{G}$.
@@
The proof is provided in Appendix C.3. Theorem 4.1 demonstrates that under bounded estimation error $\Delta$ and perturbation $\Gamma$, the most efficient PNE in $\mathcal{G}$ can be uniquely identified.
@@ 4.4.2 Performance Analysis
We now analyze the two metrics introduced in Section 4.2. Following (Bistritz and Leshem, 2018, 2021), the Learning PNE Subphase defines a Markov process $\mathcal{M}(\epsilon;\mathcal{G}^{\prime})$ on the states of all players $Z=\{(m_{j},a_{j},u_{j})\}_{j\in[N]}$ based on the parameter $\epsilon$ and the game $\mathcal{G}^{\prime}$. Based on this observation, we could obtain the following results. Note that $c_{1},c_{2},c_{3},\eta$ are hyper-parameters of our algorithm as shown in Figure 1.
@@
$\mathbb{E}[\operatorname{Reg}_{j}(T)]$ $\leq O\left(\mu_{\max}\left(c_{1}+c_{2}\log^{1+\eta}T+c_{3}\log T\right)\right),$ (14)
@@
The proof of Theorem 4.2 is detailed in Appendix C.4. The theorem establishes that our algorithm converges to the most efficient PNE in the long term. It further characterizes the regret for each player as $O(\log^{1+\eta}T)$ for any $\eta>0$, under sufficient conditions. This aligns with existing results in (Bistritz and Leshem, 2021). Additionally, as $\eta\to 0$, our bounds reduce to $O(\log T)$, consistent with standard regret bounds in the multi-armed bandit literature (Boursier and Perchet, 2022; Slivkins et al., 2019).
@@ 5.3 Analysis
We conducted experiments across 50 unique simulations by repeatedly sampling the resources’ payoff distributions and players’ weights. Our method, alongside the baseline approaches, was evaluated by computing the average regret across all players and quantifying the number of rounds where players diverged from the most efficient PNE. The experimental results are depicted in Figure 2.
@@ 6 Conclusions
In summary, we introduce the Proportional Payoff Allocation Game (PPA-Game) as a novel framework for analyzing online content creator competition, confirming the frequent emergence of pure Nash equilibrium (PNE) through extensive simulations. Further, our study extends to the Multi-player Multi-Armed Bandit (MPMAB) framework, exploring creators’ adaptive strategies in dynamic demand environments, and validates an effective new online algorithm for content optimization, backed by theoretical and experimental analysis.
Acknowledgements
Peng Cui’s research was supported in part by National Natural Science Foundation of China (No. 62141607). Bo Li’s research was supported by the National Natural Science Foundation of China (No.72171131, 72133002); the Technology and Innovation Major Project of the Ministry of Science and Technology of China under Grants 2020AAA0108400 and 2020AAA0108403.
@@ A.1 The Non-existence of PNE in Certain Scenarios
Let’s consider a configuration with $N=7$ players and $K=4$ resources. Assign the resources’ payoffs as $\bm{\mu}=(1,0.4904,0.2447,0.1065)$. The matrix representing the weights of the players for each resource is given by Equation (15). A comprehensive enumeration of possible strategy profiles confirms the non-existence of any PNE in this particular configuration.
@@
$\bm{W}=\begin{pmatrix}0.1657&0&0.5755&0\\ 0&0.0314&0&0.5623\\ 0&0&0.4217&0.9517\\ 0&1&0.4057&0\\ 1&0&0&0\\ 0&0&1&0\\ 0&0&0&1\\ \end{pmatrix}.$ (15)
Figure 3: A showcase in which PNE does not exist.
We find an improvement cycle for players $1$, $2$, $3$, and $4$ in this example, as depicted in Figure 3. In this figure, the large circles labeled A, B, C, and D represent resources. The blue numbers between these resource circles signify the players and the sequence of their actions, while the numbers inside the larger circles indicate the respective rounds. At round $1$, player $1$ perceives that resource $B$ offers a more favorable payoff and hence shifts to resource $B$. This transition decreases the payoff for player $2$, prompting him to switch to resource $C$ by round $2$. The cycle continues, and by round $4$, player $4$ transitions from resource $D$ to $B$, which subsequently reduces the payoff for player $1$. This leads player $1$ to revert to resource $A$ by round $5$, and in turn, prompts player $2$ to shift back from resource $C$ to $B$ by round $6$, thus perpetuating the cycle.

### v2

@@ 3.3.2 Existence of PNE in Specific Scenarios
$\forall k\neq k^{\prime},\quad\frac{\mu_{k}}{\mu_{k^{\prime}}}>\frac{N_{0}}{\epsilon_{0}}\quad\text{or}\quad\frac{\mu_{k}^{\prime}}{\mu_{k}}>\frac{N_{0}}{\epsilon_{0}}.$ (2)
@@
Theorem 3.1 demonstrates scenarios with a long-tailed distribution of resource payoffs. In particular, Eq. 2 specifies that the peak payoff among resources substantially surpasses that of other resources. For instance, entertainment-related content on TikTok garners substantially more views than other topics (Statista, 2023).
Several degenerate cases also guarantee the existence of a PNE.
@@
Theorem 3.2 establishes the existence of PNE in several degenerate cases of the PPA-Game (Definition 3.2). Specifically, the partially heterogeneous scenario corresponds to cases where player weights reflect their inherent influence. In online recommender platforms, for example, certain creators—termed ”influencers”—exert greater influence across all topics due to their expertise, stature, or authority. Notably, the partially heterogeneous scenario forms a special case of the weighted congestion game (Bhawalkar et al., 2014), where PNE may not always exist, yet we prove its existence in this specific setting. The homogeneous player scenario considers cases where all players exert equal influence on every topic, aligning with the model studied in (Xu et al., 2023). Meanwhile, the homogeneous resource scenario examines settings where all resources provide equivalent payoffs, as seen in competition for processing resources in modern data centers, where each resource has the same processing power (Benblidia et al., 2021).
@@ 3.3.3 Non-uniqueness and Inefficiency of PNEs
Given the existence of such inefficient PNEs, it becomes preferable for players to converge to an optimal PNE, thereby maximizing the overall welfare.
3.3.4 Price of Anarchy
We further analyze the price of anarchy (PoA) in the PPA-Game. When a PNE exists, PoA is defined as the ratio of the maximal social welfare across all strategy profiles to the minimal social welfare among all pure Nash equilibria, i.e.,
$\mathrm{PoA}=\frac{\max_{\bm{\pi}\in[K]^{N}}W_{\mathcal{G}}(\bm{\pi})}{\min_{\bm{\pi}\in[K]^{N}:\bm{\pi}\text{ is a PNE}}W_{\mathcal{G}}(\bm{\pi})}.$ (3)
Theorem 3.3.
If a PNE exists, then
$\mathrm{PoA}\leq 1+\frac{\min\{N,K\}-1}{N}.$
Remark 3.3.
Notably, this bound on PoA remains identical to that in (Xu et al., 2023), which analyzes the homogeneous player scenario in Theorem 3.2.
@@ 4.1 Multi-Player Multi-Armed Bandits Framework
When multiple players select the same arm, the reward is allocated according to the proportional payoff rule specified in Eq. 1. Specifically, let $R_{j}(t)$ represent the reward received by player $j$ at round $t$, given by:
$R_{j}(t)=X_{k}(t)\cdot\frac{w_{j,k}}{\sum_{j^{\prime}=1}^{N}\mathbb{I}[\pi_{j^{\prime}}(t)=k]w_{j^{\prime},k}},\,\text{where}\quad k=\pi_{j}(t).$
@@ 4.2 Evaluation metrics
Notably, $\operatorname{Reg}_{j}(T)$ is a random variable and the randomness comes from the actions $\{\bm{\pi}(t)\}$, which rely on historical random observations $\{X_{k}(t)\}$.
@@ 4.3 Algorithm
Inspired by (Bistritz and Leshem, 2018, 2021; Pradelski and Young, 2012), we introduce a novel online learning algorithm designed to enable players to both learn the game’s parameters and optimize their individual rewards concurrently. The algorithm’s framework is shown in Fig. 1 and the pseudo-code is outlined in Algorithm 1.
@@ 4.3.2 Learning and Exploitation Phase
In this subphase, we adopt the framework proposed in (Pradelski and Young, 2012) to identify the most efficient Nash equilibrium. Specifically, each player maintains a status tuple $(m_{j},a_{j},u_{j})$ at each round $t$ during this phase. Here, $m_{j}$ can be one of four moods: discontent ($D$), content ($C$), watchful ($C^{-}$), and hopeful ($C^{+}$). $a_{j}$ denotes the arm currently being played by player $j$, and $u_{j}$ represents the utility player $j$ gains when playing arm $a_{j}$.
Intuitively, a player’s mood reflects their attitude toward their current strategy, similar to how individuals respond emotionally or cognitively to decisions. A discontent player, uncertain about the best choice, explores alternatives, akin to someone reconsidering their options when dissatisfied. A content player, confident in their strategy, sticks with it, much like someone satisfied with their decisions. A watchful player, cautious about their strategy’s effectiveness, may shift to a discontent mood if doubts arise. Finally, a hopeful player, believing in the potential for better returns, persists with their current strategy, similar to someone holding on to a risky investment in hopes of future gains.
Following the principles underlying mood design, at each round $t$, players select arms according to Rule 4.1, based on their mood from the previous round.
@@
The chosen arm $\pi_{j}(t)$ is chosen based on the mood in the last round.
@@
Rule 4.1 encapsulates the different player moods. Discontent players, uncertain about the best strategy, explore the arm space uniformly, seeking alternatives. Content players, confident in their strategy, mostly stick with it but occasionally explore other options out of curiosity. Watchful and hopeful players, still assessing their strategy’s effectiveness, continue with their current choice, adopting a cautious or optimistic wait-and-see approach.
@@
Let $10<\epsilon<$ 1 be a constant. Define the functions $F(u)$ and $G(u)$ as follows:
@@
    $(m_{j},a_{j},u_{j})\leftarrow\begin{cases}(C^{+},a_{j},u_{j})&\text{if }u^{\prime}_{j}>u_{j},\\ (C,a_{j},u_{j})&\text{if }u^{\prime}_{j}=u_{j},\\ (C^{-},a_{j},u_{j})&\text{if }u^{\prime}_{j}<u_{j}.\end{cases}$
@@
    $(m_{j},a_{j},u_{j})\leftarrow\begin{cases}(C^{+},a_{j},u_{j})&\text{if }u^{\prime}_{j}>u_{j},\\ (C,a_{j},u_{j})&\text{if }u^{\prime}_{j}=u_{j},\\ (D,a_{j},u_{j})&\text{if }u^{\prime}_{j}<u_{j}.\end{cases}$
@@
    $(m_{j},a_{j},u_{j})\leftarrow\begin{cases}(C,a_{j},u^{\prime}_{j})&\text{if }u^{\prime}_{j}>u_{j},\\ (C,a_{j},u_{j})&\text{if }u^{\prime}_{j}=u_{j},\\ (C^{-},a_{j},u_{j})&\text{if }u^{\prime}_{j}<u_{j}.\end{cases}$
@@
Rule 4.2 reflects the intuition behind player moods. Discontent players transition to content with some probability if their chosen strategy yields high utility. The design of $F(u)$ and $\epsilon$ ensures that $\epsilon^{F(u_{j}^{\prime})}$ increases with $u_{j}^{\prime}$, meaning higher utility raises the likelihood of selecting that strategy.
Content players update their status based on the observed utility for the current round unless they have already explored other strategies, in which case they transition to a new strategy probabilistically based on utility. The design of $G(u)$ and $\epsilon$ ensures that $\epsilon^{G(u_{j}^{\prime}-u_{j})}$ increases with $u_{j}^{\prime}$, meaning higher utility raises the probability of switching strategies. Unlike discontent players, content players compare new utility to their previous utility, as captured by the $u_{j}^{\prime}-u_{j}$ term in the probability.
Watchful and hopeful players adjust their status based on observed utility. A watchful player becomes discontent if their strategy has lower utility for two consecutive rounds, while a hopeful player, upon observing utility improvement for two consecutive rounds, remains with the current strategy and transitions to content.
Consistent with existing work (Pradelski and Young, 2012; Young, 2009; Marden et al., 2014), Rules 4.1 and 4.2 follows the principles of mood design, forming a trial-and-error learning process similar to common human learning routines (Young, 2009).
Additionally, in the Learning PNE Subphase, each player maintains a counter $Q_{j,k}$ for each arm $k\in[K]$, tracking how often arm $k$ is chosen while the player is in a Content mood.
@@ 4.4.1 Quantifying the Challenge of Identifying the Most Efficient PNE
$\delta^{\text{NE}}$ $:\displaystyle=\min\big{\{}\mathcal{U}_{j}(\bm{\pi})-\mathcal{U}_{j}(\pi^{\prime},\bm{\pi}_{-j})$ : (8)
$\quad\quad\quad\,\,\,\bm{\pi}\in A^{\text{NE}},j\in[N],\pi^{\prime}\in[K],\pi^{\prime}\neq\pi_{j}\big{\}}.$
@@
Set $N=3$ and $K=2$. Here we suppose that each player has the same weights on the resources (as per Condition (2) in Theorem 3.2). When arms’ rewards are $[1,0.7]$ and players’ weights are $[1,0.8,0.4]$, we have $0.0260.052=\delta^{\text{NE}}>\delta^{\text{NoNE}}=$ 0.026. By contrast, when arms’ rewards are $[1,0.6]$ and players’ weights are $[1,2/3,4/9]$, we have $0.0680.040=\delta^{\text{NE}}<\delta^{\text{NoNE}}=$ 0.068. Details of the two cases can be found in Tables 1 and 2 in Appendix B.
Let $\Delta$ be the largest discrepancy between the estimated and true rewards $\Delta=\max_{j,k}\left|\hat{\mu}_{j,k}-\mu_{k}\right|$. We have the following theorem.
@@
Suppose $\Delta<\delta/(4K)$, $\Gamma\leq\delta/(4N)$, and the PNE exists for the game $\mathcal{G}$ in Definition 3.2. Then the PNE exists for the game $\mathcal{G}^{\prime}$ (Eq. 6). In addition, the most efficient PNE in the game $\mathcal{G}^{\prime}$ is unique with probability $1$ and the equilibrium is a most efficient PNE in the original game $\mathcal{G}$.
@@
Theorem 4.1 demonstrates that under bounded estimation error $\Delta$ and perturbation $\Gamma$, the most efficient PNE in $\mathcal{G}$ can be uniquely identified.
@@ 4.4.2 Performance Analysis
We now analyze the two metrics introduced in Section 4.2. Following (Bistritz and Leshem, 2018, 2021), the Learning PNE Subphase defines a Markov process $\mathcal{M}(\epsilon;\mathcal{G}^{\prime})$ on the states of all players $Z=\{(m_{j},a_{j},u_{j})\}_{j\in[N]}$ based on the parameter $\epsilon$ and the game $\mathcal{G}^{\prime}$. Based on this observation, we could obtain the following results. Note that $c_{1},c_{2},c_{3},\eta$ are hyper-parameters of our algorithm.
@@
$\mathbb{E}[\operatorname{Reg}_{j}(T)]$ $\leq O\left(\mu_{\max}\left(c_{1}+c_{2}\log^{1+\eta}T+c_{3}\log T\right)\right),$
@@
The theorem establishes that our algorithm converges to the most efficient PNE in the long term. It further characterizes the regret for each player as $O(\log^{1+\eta}T)$ for any $\eta>0$, under sufficient conditions. This aligns with existing results in (Bistritz and Leshem, 2021). Additionally, as $\eta\to 0$, our bounds reduce to $O(\log T)$, consistent with standard regret bounds in the multi-armed bandit literature (Boursier and Perchet, 2022; Slivkins et al., 2019).
In addition, we emphasize that the minimum value of $T$ required to guarantee the bounds in Theorem 4.2 implicitly depends on all game parameters and hyperparameters, as is also the case in (Bistritz and Leshem, 2021). Notably, the inequalities in Theorem 4.2 do not imply that smaller values of $c_{1}$, $c_{2}$, or $c_{3}$ necessarily lead to better performance. This dependence motivates our empirical tuning of these parameters in the experiments (Section 5).
@@ 5.3 Analysis
We conducted experiments across 50 unique simulations, repeatedly sampling the resources’ payoff distributions and players’ weights. Our method, along with baseline approaches, was assessed by measuring the average regret (given in Eq. 4) across all players and tracking the number of rounds in which players deviated from the most efficient PNE (given in Eq. 5). The experimental results are presented in Fig. 2.
@@ 6 Conclusions
In summary, we introduce the PPA-Game as a new framework for analyzing online content creator competition, investigating the properties of PNE. We extend this to the Multi-player Multi-Armed Bandit framework, studying creators’ adaptive strategies in dynamic demand settings. Additionally, we propose and validate a new online algorithm for content optimization, backed by theoretical guarantees and empirical evidence, demonstrating its effectiveness in enhancing long-term creator performance.
@@ A.1 The Non-existence of PNE in Certain Scenarios
Let’s consider a configuration with $N=7$ players and $K=4$ resources. Assign the resources’ payoffs as $\bm{\mu}=(1,0.4904,0.2447,0.1065)$. The matrix representing the weights of the players for each resource is given by Equation Fig. 3. A comprehensive enumeration of possible strategy profiles confirms the non-existence of any PNE in this particular configuration.
@@
$\bm{W}=\begin{pmatrix}0.1657&0&0.5755&0\\ 0&0.0314&0&0.5623\\ 0&0&0.4217&0.9517\\ 0&1&0.4057&0\\ 1&0&0&0\\ 0&0&1&0\\ 0&0&0&1\\ \end{pmatrix}.$
We find an improvement cycle for players $1$, $2$, $3$, and $4$ in this example, as depicted in Fig. 3. In this figure, the large circles labeled A, B, C, and D represent resources. The blue numbers between these resource circles signify the players and the sequence of their actions, while the numbers inside the larger circles indicate the respective rounds. At round $1$, player $1$ perceives that resource $B$ offers a more favorable payoff and hence shifts to resource $B$. This transition decreases the payoff for player $2$, prompting him to switch to resource $C$ by round $2$. The cycle continues, and by round $4$, player $4$ transitions from resource $D$ to $B$, which subsequently reduces the payoff for player $1$. This leads player $1$ to revert to resource $A$ by round $5$, and in turn, prompts player $2$ to shift back from resource $C$ to $B$ by round $6$, thus perpetuating the cycle.